OpenStudy (anonymous):

can someone help pls. The table shows the outputs y for different inputs x: Input (x) 3 7 11 15 Output (y) 4 6 8 10 Part A: Do the data in this table represent a function? Justify your answer. (3 points) Part B: Compare the data in the table with the relation f(x) = 5x – 21. Which relation has a greater value when x = 11? (2 points) Part C: Using the relation in Part B, what is the value of x if f(x) = 99? (5 points)

OpenStudy (anonymous):

A. Yes because for every input, you have an output (?)

OpenStudy (anonymous):

Because each input is related to exactly one output.

OpenStudy (anonymous):

In Part B, you need to find the value of x first in 5x-21, do you know how to do that? If you do, you'll just compare it and conclude your answer!

OpenStudy (anonymous):

This is what i got for Part B: f(11)=5(11)-21 @ienarancillo

OpenStudy (anonymous):

Ohh, I had a different way of understanding it :((

OpenStudy (anonymous):

from what I understand, there's this f(x) = 5x - 21 right? So, I guess you need to find x first And that is ..?

OpenStudy (anonymous):

honestly i have no idea im so confused.. @ienarancillo

OpenStudy (anonymous):

just find x in 5x-21 :)

OpenStudy (anonymous):

so its 34? because i plugged in the 11 for x

OpenStudy (anonymous):

yeah, meaning that the function in part b has a greater value

OpenStudy (anonymous):

as for C, plug in 99 for x

OpenStudy (anonymous):

uhm wait a sec, \[f(x)=y=5x-21\]

OpenStudy (anonymous):

\[y=\frac{ 5x }{ 5 }=\frac{ 21 }{ 5 }\] \[y=f(x)=\frac{ 21 }{ 5 } \approx 4.2\]

OpenStudy (anonymous):

but looking at the table, when x = 11 , y = 8

OpenStudy (anonymous):

so comparing y = 4.2 y = 8 I think you know it form here :)

OpenStudy (anonymous):

but please correct me if I'm wrong bc this is how I understood the problem

OpenStudy (anonymous):

im still confused :(

OpenStudy (anonymous):

let's start from the top

OpenStudy (anonymous):

oso we're given f(x)=5x-21 right?

OpenStudy (anonymous):

^ from that given equation, do you know how to find x?

OpenStudy (anonymous):

* guys please correct me if I'm wrong! :)

OpenStudy (anonymous):

yeah i plugged in the 11 for x

OpenStudy (anonymous):

Noo, if you plug in x, you'll get y. just dont mind the x=11 first, that's a different function

OpenStudy (anonymous):

like y = 5(11)-21 You get y here! Not x! :)

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

yeaah

OpenStudy (anonymous):

sooo transposing .. we get \[5x = 21\]

OpenStudy (anonymous):

and then, we want to find x so divide both sides by 5 \[\frac{ 5x }{ 5 } = \frac{ 21 }{ 5 }\]

OpenStudy (anonymous):

so far so good?